# Self-supervised pre-training with diffusion model for few-shot landmark detection in x-ray images

### v1

@@ 6 Conclusions
Despite the widespread application of deep neural networks for medical image analysis, data scarcity in images and annotations represents a fundamental challenge for this domain. In this research, we propose and evaluate the impact of a pre-training approach based on self-supervision with DDPM for label-efficient landmark detection tasks in x-ray images, benchmarking it against the typically adopted supervised ImageNet pre-training and self-supervised MoCoV3, SimCLRV2, and DINO. Specifically, given an x-ray image dataset, we leverage the training set in a self-supervised fashion for pre-training a DDPM based on a Unet, further fine-tuning it on the available annotated training images. Our results on three popular benchmark datasets show that the proposed method outperforms ImageNet supervised pre-training and MoCoV3, SimCLRV2, DINO self-supervised pre-training in the context of few annotations available for the landmark detection task. Moreover, such improvement seems to be preserved when pre-training on a different in-domain dataset and fine-tuning on the downstream one, simulating the interesting case where both annotations and images are limited for the downstream task. We believe that the proposed method showcases the potential of self-supervised diffusion-based pre-training for the landmark detection task, potentially opening new research directions to effectively address data scarcity challenges in this domain.

### v3

@@ 6 Conclusions
Our research introduces a novel application of DDPMs for anatomical landmark detection in X-ray images, addressing the persistent challenge of limited annotated data in medical image analysis. This work bridges the gap between the established potential of diffusion models in medical imaging and the critical need for label-efficient methods in landmark localization.
We demonstrate that our DDPM-based self-supervised pre-training method consistently outperforms supervised ImageNet pre-training, traditional state-of-the-art self-supervised approaches (MoCoV3, SimCLRV2, DINO), and the YOLO framework in few-shot learning scenarios across Chest, Cephalometric, and Hand X-ray datasets. The performance gains are particularly significant in low-data regimes, with substantial MRE reductions and SDR improvements. Notably, our method maintains its effectiveness when pre-trained on one in-domain dataset and fine-tuned on smaller, distinct datasets, simulating real-world scenarios with limited data availability. This robustness suggests broad applicability across various medical imaging tasks and datasets.
Our work not only addresses the critical issue of data scarcity but also contributes to the ongoing discussion on effective pre-training strategies in medical image analysis. By showing the efficacy of DDPMs in this context, we aims to open new avenues for research into efficient and robust landmark detection methods, potentially accelerating progress in automated diagnosis systems.